# Primordial neutrinos fade to gray: constraints from cosmological observables

### v1

@@ Abstract
We investigate the effect of potentially large distortions of the relic neutrino spectra on cosmological observables. To that end, we consider a phenomenological model of “gray” spectral distributions, described by a single parameter which generalizes the traditional $y$-distortions to possibly large negative values. Implementing these distortions in the primordial nucleosynthesis code PRIMAT, we can constrain the distortion parameter along with the presence of extra radiation, exploiting the complementarity of big bang nucleosynthesis and cosmic microwave background measurements to disentangle gravitational and non-thermal effects. These constraints rule out a distortion where more than $\sim 1/2$ of the neutrinos’ energy density is replaced by dark radiation. Nonetheless, we find that large distortions, accompanied with extra radiation, are allowed—and even slightly preferred in some cases—by current cosmological observations. As this scenario would require substantial modifications to the physics of neutrino decoupling in the early Universe, these observational constraints call for a renewed attention on the possibility of large deviations from the standard cosmological model in the neutrino sector.
@@ I Introduction
The effect of neutrinos can be seen in various cosmological observables Lesgourgues et al. (2013), such as the Cosmic Microwave Background (CMB). It is primarily sensitive to the energy density of the cosmic neutrino background, allowing to measure the effective number of relativistic species $N_{\mathrm{eff}}$, consistently with the standard prediction of $3.044$ Froustey et al. (2020); Bennett et al. (2021); Drewes et al. (2024). A complementary probe is given by spectroscopic measurements of the primordial abundances, since neutrinos set the conditions under which BBN takes place Grohs and Fuller (2023). The agreement between theory (with publicly available codes like PRIMAT Pitrou et al. (2018), PArthENoPE Gariazzo et al. (2022) or PRyMordial Burns et al. (2024)) and observations for deuterium and helium abundances is a major success of the standard cosmological model, slightly obscured by the so-called lithium problem Fields (2011). While these two probes are sensitive to the epoch when neutrinos were ultrarelativistic, the growth of large-scale structures is affected by the neutrino masses, which are known to be nonzero because of the phenomenon of neutrino oscillations Kajita (2016); McDonald (2016). However, this cosmological effect has not been seen yet. Quite the opposite — the recent results by DESI Adame et al. (2025); Abdul Karim et al. (2025) do not show preference for massive neutrinos, and if one allows for negative effective masses, there is a $\sim 3\sigma$ tension between the measurements and the minimum value compatible with neutrino flavor oscillations Craig et al. (2024).
@@
This growing tension calls for the study of non-standard features in the cosmic neutrino spectrum. Moreover, although early Universe measurements (CMB, BBN) are consistent with the standard scenario of neutrino decoupling, it is worth emphasizing that they are only indirect probes of the early cosmic history. Specifically, highly non-standard scenarios which are still compatible with measurements simply cannot be excluded Cuoco et al. (2005); Alvey et al. (2022). In this article, we thus focus on non-standard neutrino spectral distortions.
@@
In the standard scenario, percent-level non-thermal distortions are present because of the out-of-equilibrium decoupling process Dolgov et al. (1997); Mangano et al. (2005); Grohs et al. (2016), but their effects are small on cosmological observables (see e.g., Froustey and Pitrou (2020)). A popular scenario assumes nonzero neutrino/antineutrino asymmetries, which can be associated to nonzero chemical potentials. Such a change modifies in particular the neutron-to-proton ratio at the onset of BBN, allowing to constrain these asymmetries Simha and Steigman (2008); Shimon et al. (2010); Castorina et al. (2012); Barenboim et al. (2017); Oldengott and Schwarz (2017); Escudero et al. (2023); Froustey and Pitrou (2024); Lattanzi and Moretti (2024); Domcke et al. (2025). Here, we consider potentially large deviations from the standard Fermi-Dirac spectrum, identical for neutrinos and antineutrinos, starting from $y$-type distortions similar to the ones introduced for the neutrino spectrum in Barenboim et al. (2024). We generalize them to describe both an increase or decrease of the neutrino energy density, and allow for the presence of additional dark radiation. Specifically, we consider a so-called “gray” distribution, a type of distribution usually discussed in the context of CMB (see, e.g., Stebbins (2007); Pitrou and Stebbins (2014)).
@@
This paper is structured as follows: in section II, we introduce our model for a distorted cosmic neutrino spectrum, and study its theoretical implications for cosmological observables, with special emphasis on BBN. In section III, we employ the BBN code PRIMAT to quantitatively describe the effects of this distorted spectrum. Then, in section IV, we present constraints on our model from cosmological observations. Finally, we discuss the implications of those constraints and conclude in section V. Throughout this paper, we use natural units where $\hbar=c=k_{B}=1$.
@@ II.1.1 $y$ and $\mu$-type distortions
In the CMB, nonthermal features in the distribution function of background photons can be divided into two distinct types: $\mu$-type and $y$-type distortions. These distortions, which are well known, well studied, and bounded to be very small in the CMB spectrum, can arise from inefficient (efficient) Compton scattering with electrons for the $y$ ($\mu$) type. In the case of the cosmic neutrino background, the same types of distortions can be used. Furthermore, $\mu$-type distortions can be attributed to the chemical potentials of (anti)neutrinos, which emerge if the lepton asymmetry of the Universe is significant, a possibility allowed by current experimental data Simha and Steigman (2008); Shimon et al. (2010); Castorina et al. (2012); Barenboim et al. (2017); Oldengott and Schwarz (2017); Escudero et al. (2023); Froustey and Pitrou (2024); Lattanzi and Moretti (2024); Domcke et al. (2025). In this case, however, these “distortions” should not be interpreted as such, since a chemical potential is a thermal albeit non-standard feature of the neutrino distribution function.
@@
where $\hat{\rho}_{\nu}$ (resp. $\hat{\rho}_{\nu}+\delta\rho_{\nu}$) is the energy density calculated with $\hat{f}_{\nu}$ (resp. $f_{\nu}(x,y_{\mathrm{SZ}})$). Another little-known feature is that the $y_{\mathrm{SZ}}$ parameter must be positive. This limitation can be understood by seeing the spectrum as a sum of blackbodies [^1] Although for photons the “blackbody” spectrum is the Planck distribution, we can also use this term for neutrinos, implying a Fermi-Dirac distribution. with different temperatures, and characterize the distribution of the temperatures by its moments in temperature space. Stebbins Stebbins (2007) has shown that the superposition of blackbodies is better characterized as a distribution in the variable $\log(T)$ and not $T$. By doing so, it can be easily demonstrated Pitrou and Stebbins (2014) that the average of $\log(T)$ corresponds to the average of the temperatures, while the variance of the distribution, obviously a positive quantity, can be mapped directly to what is usually called a $y_{\mathrm{SZ}}$-distortion. More directly, one can also see that the function (1) is asymptotically equivalent, for $x\to\infty$, to $y_{\mathrm{SZ}}x^{2}e^{-x}$, which has to be positive for $f_{\nu}$ to be a distribution function. Extensions and higher order corrections, normally called down-Comptonization (see, e.g., Miao et al. (2021)), do exist in the literature, however they break down at relatively low energy depletion.
@@ II.1.2 Generalized distortions
$f_{\nu}(x)=\int_{0}^{\infty}p(t)\,\hat{f}_{\nu}\left(\frac{x}{t}\right)\mathrm{d}t$ (5)
where $p(t)$ is a distribution function of temperatures $t$ (in units of the reference temperature $T$). If $\int_{0}^{\infty}p(t)\,\mathrm{d}t\neq 1$ the distribution is called “gray”, and the grayness parameter Stebbins (2007) is defined by $1-g\equiv\int_{0}^{\infty}p(t)\,\mathrm{d}t$. A full description of gray distributions and a complete generalization of the $y_{\mathrm{SZ}}$ distortion is beyond the scope of this work and will be given elsewhere.
In this framework, a distribution of relative temperatures which preserves the number density while modifying the energy density can be achieved for $y>0$ with the distribution
@@ II.1.3 Distortions and incomplete neutrino decoupling
Specifically, we need to distinguish the temperature of the electromagnetic plasma ($T_{\gamma}$), shared by photons, electrons, positrons and baryons, and the temperature $T_{\nu}$ neutrinos would have without distortions. In our approximation, neutrinos have decoupled such that $T_{\nu}\propto a^{-1}$, with $a$ the scale factor, while $T_{\gamma}$ is modified by $e^{+}e^{-}$ annihilations, which leads to the standard ratio $T_{\nu}/T_{\gamma}\to(4/11)^{1/3}$ after annihilations (if one neglects QED corrections to the plasma thermodynamics). The distorted distribution $f_{\nu}(x,y_{g})$ is defined with respect to this temperature $T_{\nu}$, such that $x=E/T_{\nu}$ in Eq. (8), with $E$ the neutrino energy.
@@ II.2.1 Energy density and $N_{\mathrm{eff}}$ parameter
After electron-positron annihilations (i.e., for $T_{\gamma}\ll m_{e}\simeq 0.511\,\mathrm{MeV}$), the energy density of ultrarelativistic species can be written
@@
More specifically, we define $\Delta N$ such that the case $\Delta N=0$ (with $y_{g}=0$ as well) corresponds to $N_{\mathrm{eff}}=N_{\mathrm{eff}}^{\mathrm{std}}\simeq 3.010$. This value corresponds to the standard prediction in the instantaneous decoupling scenario (see discussion above), including QED corrections to the plasma thermodynamics (see, e.g., Table 1 in Bennett et al. (2020) or Table I in Froustey and Pitrou (2020)). Making this instantaneous decoupling approximation means that our $N_{\mathrm{eff}}$ predictions may have an error of the order of 1 %, since the “true” standard value is $N_{\mathrm{eff}}^{\mathrm{std}}=3.044$ Froustey et al. (2020); Bennett et al. (2021); Drewes et al. (2024). This error is still well below current experimental uncertainties on $N_{\mathrm{eff}}$. In our analysis, the observational constraint on $N_{\mathrm{eff}}$, which is a constraint on $\rho_{\mathrm{rad}}$, allows us to constrain simultaneously the distortion parameter $y_{g}$ and $\Delta N$, since $N_{\mathrm{eff}}^{\mathrm{obs}}=N_{\mathrm{eff}}^{\mathrm{std}}(1+y_{g})+\Delta N$, thanks to Eqs. (9), (10) and (11).
@@
Finally, we note that we will also allow the $\Delta N$ parameter to be negative. Even though this is incompatible with the interpretation of $\Delta N$ as additional dark radiation, there are a range of physical scenarios that can decrease $N_{\mathrm{eff}}$, which would correspond to $\Delta N<0$, see e.g., Steigman (2013); Chluba et al. (2020); Sanchis et al. (2025); Cadamuro and Redondo (2012); Fuller et al. (2011); Boyarsky et al. (2021); Rasmussen et al. (2022); Ovchynnikov and Syvolap (2025); Akita et al. (2025). Current CMB experiments do not exclude $N_{\mathrm{eff}}<3$, see for instance the recent ACT results Louis et al. (2025); Calabrese et al. (2025).
@@ II.2.2 Neutrinos and BBN
At the onset of BBN, almost all free neutrons are fused into helium-4, with small amounts of deuterium and traces of heavier elements (e.g., lithium). The final helium abundance is then directly set by the amount of free neutrons “available” at BBN. This number is usually considered via the neutron-to-proton ratio $n_{n}/n_{p}$, which is set by the following weak reactions:

### v2

@@ Abstract
We investigate the effect of potentially large distortions of the relic neutrino spectra on cosmological observables. To that end, we consider a phenomenological model of “gray” spectral distributions, described by a single parameter, which generalizes the traditional $y$-distortions to possibly large negative values. Implementing these distortions in the primordial nucleosynthesis code PRIMAT, we can constrain the distortion parameter along with the presence of extra radiation, exploiting the complementarity of big bang nucleosynthesis and cosmic microwave background measurements to disentangle gravitational and non-thermal effects. These constraints rule out a distortion where more than $\sim 1/2$ of the neutrinos’ energy density is replaced by dark radiation. Nonetheless, we find that large distortions, accompanied with extra radiation, are allowed—and even slightly preferred in some cases—by current cosmological observations. As this scenario would require substantial modifications to the physics of neutrino decoupling in the early Universe, these observational constraints call for a renewed attention on the possibility of large deviations from the standard cosmological model in the neutrino sector.
@@ I Introduction
The effect of neutrinos can be seen in various cosmological observables Lesgourgues et al. (2013), such as the Cosmic Microwave Background (CMB). It is primarily sensitive to the energy density of the cosmic neutrino background, allowing one to measure the effective number of relativistic species $N_{\mathrm{eff}}$, consistently with the standard prediction of $3.044$ Froustey et al. (2020); Bennett et al. (2021); Drewes et al. (2024). A complementary probe is given by spectroscopic measurements of the primordial abundances, since neutrinos set the conditions under which BBN takes place Grohs and Fuller (2023). The agreement between theory (with publicly available codes like PRIMAT Pitrou et al. (2018), PArthENoPE Gariazzo et al. (2022) or PRyMordial Burns et al. (2024)) and observations for deuterium and helium abundances is a major success of the standard cosmological model, slightly obscured by the so-called lithium problem Fields (2011). While these two probes are sensitive to the epoch when neutrinos were ultrarelativistic, the growth of large-scale structures is affected by the neutrino masses, which are known to be nonzero because of the phenomenon of neutrino oscillations Kajita (2016); McDonald (2016). However, this cosmological effect has not been seen yet. Quite the opposite — the recent results by DESI Adame et al. (2025); Abdul Karim et al. (2025) do not show preference for massive neutrinos, and if one allows for negative effective masses, there is a $\sim 3\sigma$ tension between the cosmological measurements and the minimum value compatible with neutrino flavor oscillations Craig et al. (2024); Jiang et al. (2025).
@@
This growing tension calls for the study of non-standard features in the cosmic neutrino spectrum. Moreover, although early Universe measurements (CMB, BBN) are consistent with the standard scenario of neutrino decoupling, it is worth emphasizing that they are only indirect probes of the early cosmic history. Specifically, highly non-standard scenarios that are still compatible with measurements simply cannot be excluded Cuoco et al. (2005); Alvey et al. (2022). In this article, we thus focus on non-standard neutrino spectral distortions.
@@
In the standard scenario, percent-level non-thermal distortions are present because of the out-of-equilibrium decoupling process Dolgov et al. (1997); Mangano et al. (2005); Grohs et al. (2016), but their effects are small on cosmological observables (see e.g., Froustey and Pitrou (2020)). A popular scenario assumes nonzero neutrino/antineutrino asymmetries, which can be associated to nonzero chemical potentials. Such a change modifies in particular the neutron-to-proton ratio at the onset of BBN, allowing one to constrain these asymmetries Simha and Steigman (2008); Shimon et al. (2010); Castorina et al. (2012); Barenboim et al. (2017); Oldengott and Schwarz (2017); Escudero et al. (2023); Froustey and Pitrou (2024); Lattanzi and Moretti (2024); Domcke et al. (2025). Here, we consider potentially large deviations from the standard Fermi-Dirac spectrum, identical for neutrinos and antineutrinos, starting from $y$-type distortions similar to the ones introduced for the neutrino spectrum in Barenboim et al. (2024). We generalize them to describe both an increase or decrease of the neutrino energy density, and allow for the presence of additional dark radiation. Specifically, we consider a so-called “gray” distribution, a type of distribution usually discussed in the context of CMB (see, e.g., Stebbins (2007); Pitrou and Stebbins (2014)).
@@
This paper is structured as follows: in Sec. II, we introduce our model for a distorted cosmic neutrino spectrum, and study its theoretical implications for cosmological observables, with special emphasis on BBN. In Sec. III, we employ the BBN code PRIMAT to quantitatively describe the effects of this distorted spectrum. Then, in Sec. IV, we present constraints on our model from cosmological observations. Finally, we discuss the implications of those constraints and conclude in Sec. V. Throughout this paper, we use natural units where $\hbar=c=k_{B}=1$.
@@ II.1.1 $y$ and $\mu$-type distortions
In the CMB, nonthermal features in the distribution function of background photons can be divided into two distinct types: $\mu$-type and $y$-type distortions. These distortions, which are well known, well studied, and bounded to be very small in the CMB spectrum, can arise from inefficient (efficient) Compton scattering with electrons for the $y$ ($\mu$) type. In the case of the cosmic neutrino background, the same types of distortions can be used. Furthermore, $\mu$-type distortions can be attributed to the chemical potentials of (anti)neutrinos, which emerge if the lepton asymmetry of the Universe is significant, a possibility allowed by current experimental data Simha and Steigman (2008); Shimon et al. (2010); Castorina et al. (2012); Barenboim et al. (2017); Oldengott and Schwarz (2017); Escudero et al. (2023); Froustey and Pitrou (2024); Lattanzi and Moretti (2024); Domcke et al. (2025). In this case, however, these “distortions” should not be interpreted as such, since a chemical potential is a thermal albeit non-standard feature of the neutrino distribution function. We also note that low reheating temperature scenarios can lead to $\mu$-type-like distortions, and the impact on cosmological observables allows one, in turn, to constrain these scenarios in the same spirit as this work Kawasaki et al. (1999, 2000); de Salas et al. (2015); Hasegawa et al. (2019); Barbieri et al. (2025).
@@
where $\hat{\rho}_{\nu}$ (resp. $\hat{\rho}_{\nu}+\delta\rho_{\nu}$) is the energy density calculated with $\hat{f}_{\nu}$ [resp. $f_{\nu}(x,y_{\mathrm{SZ}})$]. Another little-known feature is that the $y_{\mathrm{SZ}}$ parameter must be positive. This limitation can be understood by seeing the spectrum as a sum of blackbodies [^1] Although for photons the “blackbody” spectrum is the Planck distribution, we can also use this term for neutrinos, implying a Fermi-Dirac distribution. with different temperatures, and characterizing the distribution of the temperatures by its moments in temperature space. Stebbins Stebbins (2007) has shown that the superposition of blackbodies is better characterized as a distribution in the variable $\log(T)$ and not $T$. By doing so, it can be easily demonstrated Pitrou and Stebbins (2014) that the average of $\log(T)$ corresponds to the average of the temperatures, while the variance of the distribution, obviously a positive quantity, can be mapped directly to what is usually called a $y_{\mathrm{SZ}}$-distortion. More directly, one can also see that the function (1) is asymptotically equivalent, for $x\to\infty$, to $y_{\mathrm{SZ}}x^{2}e^{-x}$, which has to be positive for $f_{\nu}$ to be a distribution function. Extensions and higher order corrections, normally called down-Comptonization (see, e.g., Miao et al. (2021)), do exist in the literature, however they break down at relatively low energy depletion.
@@ II.1.2 Generalized distortions
$f_{\nu}(x)=\int_{0}^{\infty}p(t)\,\hat{f}_{\nu}\left(\frac{x}{t}\right)\mathrm{d}t\,,$ (5)
where $p(t)$ is a distribution function of temperatures $t$ (in units of the reference temperature $T$). If $\int_{0}^{\infty}p(t)\,\mathrm{d}t\neq 1$ the distribution is called “gray,” and the grayness parameter Stebbins (2007) is defined by $1-g\equiv\int_{0}^{\infty}p(t)\,\mathrm{d}t$. A full description of gray distributions and a complete generalization of the $y_{\mathrm{SZ}}$ distortion is beyond the scope of this work and will be given elsewhere.
In this framework, a distribution of relative temperatures, which preserves the number density while modifying the energy density, can be achieved for $y>0$ with the distribution
@@ II.1.3 Distortions and incomplete neutrino decoupling
Specifically, we need to distinguish the temperature of the electromagnetic plasma ($T_{\gamma}$), shared by photons, electrons, positrons and baryons, and the temperature $T_{\nu}$ neutrinos would have without distortions. In our approximation, neutrinos have decoupled such that $T_{\nu}\propto a^{-1}$, with $a$ the scale factor, while $T_{\gamma}$ is modified by $e^{+}e^{-}$ annihilations, which leads to a ratio $T_{\gamma}/T_{\nu}\simeq 1.40$ after annihilations. The distorted distribution $f_{\nu}(x,y_{g})$ is defined with respect to this temperature $T_{\nu}$, such that $x=E/T_{\nu}$ in Eq. (8), with $E$ the neutrino energy.
@@ II.2.1 Energy density and $N_{\mathrm{eff}}$ parameter
After electron-positron annihilations (i.e., for $T_{\gamma}\ll m_{e}\simeq 0.511\,\mathrm{MeV}$), the energy density of ultrarelativistic species can be written as
@@
More specifically, we define $\Delta N$ such that the case $\Delta N=0$ (with $y_{g}=0$ as well) corresponds to $N_{\mathrm{eff}}=N_{\mathrm{eff}}^{\mathrm{std}}\simeq 3.010$. This value corresponds to the standard prediction in the instantaneous decoupling scenario (see discussion above), including QED corrections to the plasma thermodynamics [^6] We emphasize that an “instantaneous decoupling” is not incompatible with $N_{\mathrm{eff}}\neq 3$, since a value of exactly 3 requires to assume that the $e^{+}e^{-}$ were ultrarelativistic at the time of neutrino decoupling (which we assume), and that the thermodynamics of the plasma do not include QED corrections (but we include them), see notably the Conclusions of Ref. Bennett et al. (2020). (see, e.g., Table 1 in Bennett et al. (2020) or Table I in Froustey and Pitrou (2020)). Making this instantaneous decoupling approximation means that our $N_{\mathrm{eff}}$ predictions may have an error of the order of 1 %, since the “true” standard value is $N_{\mathrm{eff}}^{\mathrm{std}}=3.044$ Froustey et al. (2020); Bennett et al. (2021); Drewes et al. (2024). This error is still well below current experimental uncertainties on $N_{\mathrm{eff}}$. In our analysis, the observational constraint on $N_{\mathrm{eff}}$, which is a constraint on $\rho_{\mathrm{rad}}$, allows us to constrain simultaneously the distortion parameter $y_{g}$ and $\Delta N$, since $N_{\mathrm{eff}}^{\mathrm{obs}}=N_{\mathrm{eff}}^{\mathrm{std}}(1+y_{g})+\Delta N$, thanks to Eqs. (9), (10) and (11).
@@
Finally, we note that we will also allow the $\Delta N$ parameter to be negative. Even though this is incompatible with the interpretation of $\Delta N$ as additional dark radiation, there are a range of physical scenarios that can decrease $N_{\mathrm{eff}}$, which would correspond to $\Delta N<0$, see e.g., Steigman (2013); Chluba et al. (2020); Sanchis et al. (2025); Cadamuro and Redondo (2012); Fuller et al. (2011); Boyarsky et al. (2021); Rasmussen et al. (2022); Ovchynnikov and Syvolap (2025); Akita et al. (2025). Current CMB experiments do not exclude this possibility, see for instance the recent ACT results Louis et al. (2025); Calabrese et al. (2025), which actually prefer $N_{\mathrm{eff}}<3$.
The gravitational effects parametrized by $N_{\mathrm{eff}}$ affect various cosmological observables, most notably the CMB. As an example, Fig. 3 shows the effect of varying $y_{g}$ (keeping $\Delta N=0$ and all other cosmological parameters to their preferred values from Planck) on the CMB temperature-temperature angular power spectrum $C_{l}^{\mathrm{TT}}$, obtained with a modified version of the code CLASS Blas et al. (2011) (see Sec. IV). We also represent the data from Planck Aghanim et al. (2020a), which shows that non-zero values of $y_{g}$, for $\Delta N=0$, would be excluded by measurements. However, if $\Delta N=-N_{\mathrm{eff}}^{\mathrm{std}}y_{g}$, the combined effects cancel and $N_{\mathrm{eff}}^{\mathrm{obs}}=N_{\mathrm{eff}}^{\mathrm{std}}$. As the CMB’s dependence on $y_{g}$ and $\Delta N$ comes mostly from this gravitational effect, this causes a degeneracy between the two parameters, as seen in Fig. 4. The solid lines show the relative difference of $C_{l}^{\mathrm{TT}}$ compared to the unperturbed baseline, with $\Delta N=0$. The dashed lines show, for the same value of $y_{g}$, the results with $\Delta N\simeq-N_{\mathrm{eff}}^{\mathrm{std}}y_{g}$, with a systematically reduced deviation from standard $\Lambda$CDM compared to $\Delta N=0$. Even though there are secondary effects that cause the cancellation not to be complete (in particular, the CMB damping tail has a mild dependence on the helium abundance, which, as we will see later, has different dependencies on $y_{g}$ and $\Delta N$), the non-cancellation is too small given current experimental bounds. We thus expect a significant degeneracy between $y_{g}$ and $\Delta N$ in our analysis from CMB measurements only (see Sec. IV).
Figure 3: Effect of $y_{g}$ distortions on the CMB TT spectrum, assuming $\Delta N=0$. The figure was generated using our modified version of CLASS Blas et al. (2011), see Sec. IV for details.
Figure 4: Relative change of the CMB TT spectrum with respect to $\Lambda\mathrm{CDM}$ in presence of $y_{g}$ distortions and $\Delta N$. We show that $\Delta N$ can roughly cancel the effect of $y_{g}$ and, even though the cancellation is not exact, the difference is too small given the current experimental sensitivity.
@@ II.2.2 Neutrinos and BBN
At the onset of BBN, almost all free neutrons are fused into helium-4, with small amounts of deuterium and traces of heavier elements (e.g., lithium). The final helium abundance is then directly set by the amount of free neutrons “available” at BBN. This number is usually quantified via the neutron-to-proton ratio $n_{n}/n_{p}$, which is set by the following weak reactions: